A girl is walking in the rain at a speed of
1.2 ms-1 along a straight road. The rain falls
on her vertically at a speed of F4 ms 1 . At
what angle to the vertical does the person
hold the umbrella so that she does not get
wet ?
1
Expert's answer
2020-09-30T12:26:04-0400

"Vm= 1.2m\/s"


"Vr= 1.4m\/s"


Let "\theta" b the angle the girl is supposed to walk.


"tan\theta= \frac {Vr}{Vm}"


"tan\theta= \frac {1.4}{1.2}"


"tan\theta= 1.167"


"\theta= 49.41\degree"